A manager of a large workforce wishes to forecast the monthly employee turnover within his workforce.

Turnover data for the last several months is given below.

month turnover
1 20
2 30
3 23
4 33
5 50

The manager decides to use a three month moving average. What is the turnover forecast for month 6?

Enter the exact answer to two decimal places of accuracy. (If the answer were an even 25, enter 25.00.)
Answer #1

the necessary calculation table:-

month turnover 3 month moving average
1 20
2 30
3 23
4 33 (20+30+23) /3 = 21.33
5 50 (30+23+33)/3 = 28.67
6 (23+33+50)/3 = 35.33

the turnover forecast for month 6 = 35.33

*** if you have any doubt regarding the problem please write it in the comment box.if you are satisfied please give me a LIKE if possible...